# From Limited Labels to Open Domains:An Efficient Learning Method for Drone-view Geo-Localization

### v3

@@ IV-C Comparison with State-of-the-art Methods
IV-D Ablation Studies
Effectiveness of CDIS. As shown in Table IX, both the SIFL block and SIEL module contribute significantly to performance improvements with both 2% and 100% of paired data. When used individually, each component increases R@1 and AP, which demonstrates their individual effectiveness. Notably, their joint optimization yields further improvement, which validates the design of the CDIS. Specifically, SIFL enhances geometric invariance at different viewpoints by modeling structural priors, while SIEL improves spatial consistency by incorporating scale-aware contextual cues. These two components complement each other to enable CDIKTNet to construct a unified and robust cross-view shared feature space.
[…]
@@ IV-E Further Analysis
Hyper-parameters Analysis of the loss coefficients. We further investigate the influence of the loss coefficients $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ in Eq. (7) on the University-1652 dataset. As shown in Fig. 4, all three hyper-parameters exhibit stable trends within reasonable ranges. For $\lambda_{1}$, our method achieves consistently strong performance for $\lambda_{3}\in[0.4,0.8]$, and overly small or large values weaken the contribution of this term. For $\lambda_{2}$, the best results are obtained around $\lambda_{2}=0.10$, while both smaller and larger values lead to slight performance degradation. For $\lambda_{1}$, the performance gradually improves as the weight increases from 0.40 to 1.00, and the best results are obtained when $\lambda_{3}=1.00$. Based on these observations, the best overall performance is achieved at $\lambda_{1}=0.60$, $\lambda_{2}=0.10$, and $\lambda_{3}=1.00$, and therefore we adopt these coefficients in all subsequent experiments.
Figure 4: Performance Influence of $\lambda_{1}$$\lambda_{2}$ and $\lambda_{3}$. Influence of different $\lambda_{1}$$\lambda_{2}$ and $\lambda_{3}$ values on University-1652 in setting iii.
@@
Influence of CDIS on Convergence. To further examine the influence of CDIS on convergence, we compare the training behavior of the baseline and baseline with CDIS on the University-1652 dataset in setting i. As shown in Fig. 6(a), it presents the loss curves and their statistical values. The baseline and baseline with CDIS follow similar descent patterns and show stable convergence regions. The baseline with CDIS begins with a slightly higher loss but reaches a lower final value, which demonstrates that the optimization process remains stable and that CDIS guides the model toward a better solution. Fig. 6(b) reports the performance and the training time. CDIS improves the performance in both the drone $\rightarrow$ satellite and satellite $\rightarrow$ drone scenarios, and the additional training cost is small at about 35 seconds. These observations show that CDIS strengthens structural and spatial invariance without causing slower convergence and adds only a minor computational cost while achieves better performance.
[…]
As shown in Table X, in setting i, CDIKTNet surpasses most supervised methods despite the complexity of the dataset. In setting ii, Although sacrifices a small portion of paired data but greatly exceeds the performance of EM-CVGL. Moreover, in setting iii, CDIKTNet is initialized through cross-domain transfer using the University-1652 dataset. It still outperforms EM-CVGL. However, in settings iiand iii, CDIKTNet performs suboptimally compared to current supervised methods. The core challenge lies in the inherent confusion within the DenseUAV dataset, particularly the issues of cross-view misalignment and spatio-temporal variations. These challenges become more pronounced when paired data is limited during the initialization phase. Additionally, cross-domain initialization requires the model to possess a certain degree of generalization ability from the source domain to the target domain for optimal initialization. This limitation highlights the importance of establishing strong prior knowledge during the initialization phase, which will significantly enhance cross-view alignment capabilities. In future work, we will explore methods that do not rely on such priors.
@@
i ConvNext-Tiny[DenseUAV] 100 60.23 81.94 - - - - - -
Sample4Geo[deuser2023sample4geo] 100 80.57 96.53 80.57 96.53 86.87 98.71 91.38 99.74
DAC[xia2024enhancing] 100 84.47 96.53 80.92 97.04 85.32 98.71 85.46 99.10
@@ IV-F Visualization
As shown in Fig. 7, to further demonstrate the effectiveness of CDIKTNet, we visualize the similarity distribution using t-SNE [van2008visualizing] and feature space projection.
Similarity Distribution Visualization. As shown in Fig. 7(a–b), Sample4Geo and DAC achieve overlap areas of 0.1232 and 0.1124 under limited supervision, due to their inability to exploit unpaired data. Fig. 7(c) shows EM-CVGL [li2024learning] achieves an overlap of 0.1158. Although it performs slightly better, feature confusion remains an issue. In contrast, CDIKTNet in Fig. 7(d) achieves a lower overlap of 0.0661 with only 2% of paired data. This confirms its strong ability to align and separate cross-view features under limited supervision. Notably, in Fig. 7(e), CDIKTNet is trained in the cross-domain initialization setting described in Subsection IV-C. It still achieves an overlap of 0.0808 to outperform all methods. This shows its effectiveness in exploiting unpaired data and transferring knowledge from an external source domain.
Feature Space Visualization. In the feature space, features from the same location should cluster, while different locations should be well separated. As shown in Fig. 7(f–h), the compared methods show noticeable cross-view confusion (red circles) and dispersed feature distributions. In contrast, CDIKTNet (Fig. 7(i–j)) produces more compact and structured feature distributions in both scenarios. Although the 0% of paired data in cross-domain initialization setting shows minor confusion, the features remain more compact and more evenly aligned. This presents the ability of CDIKTNet in learning cross-view information.
Figure 7: Similarity Distribution and Feature Space Visualization on University-1652 . (a–e) show the similarity distribution visualization. The red-shaded area indicates the overlapping region between the positive and negative distributions. (f-j) show the distribution of feature embeddings in the 2D shared feature space, where circles and triangles in different colors denote drone and satellite view. A total of 20 localizations are randomly selected from the test set.
Attention Heatmap Visualization. As a further demonstration of the effectiveness of SIEL, we visualize its attention heat maps in Fig. 8. Despite the large spatial differences between satellite and drone images and the feature distribution shifts caused by varying drone viewpoints, SIEL consistently highlights semantically meaningful regions in both views.

### v4

@@ IV-C Comparison with State-of-the-art Methods
IV-D Ablation Studies
Effectiveness of CDIS. As shown in Table IX, both the SIFL block and SIEL module contribute significantly to performance improvements with both 2% and 100% of paired data. When used individually, each component increases R@1 and AP, which demonstrates their individual effectiveness. Notably, their joint optimization yields further improvement, which validates the design of CDIS. Specifically, SIFL enhances geometric invariance at different viewpoints by modeling structural priors, while SIEL improves spatial invariance by incorporating scale-aware contextual cues. These two components complement each other to enable CDIKTNet to construct a unified and robust cross-view shared feature space.
[…]
@@ IV-E Further Analysis
Hyper-parameters Analysis of Loss Coefficients. We further investigate the influence of the loss coefficients $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ in Eq.  (7) on the University-1652 dataset. As shown in Fig. 4, the performance influence of all three hyper-parameters exhibit bounded fluctuations within reasonable ranges. It achieves consistently robust performance for $\lambda_{1}\in[0.4,0.8]$, and values outside this range diminish the effectiveness. Then the peak performance occurs near $\lambda_{2}=0.10$, with marginal degradation observed at both lower and higher values. Furthermore, it exhibits a gradual improvement as its weight increases from 0.40 to 1.00 for $\lambda_{3}$, in which the best results are obtained when $\lambda_{3}=1.00$. Synthesizing these observations, the highest overall performance is attained at $\lambda_{1}=0.60$, $\lambda_{2}=0.10$, and $\lambda_{3}=1.00$. Consequently, these coefficient values are adopted for all subsequent experiments.
Figure 4: Performance Influence of $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$. Influence of different $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ values on University-1652 in setting iii.
@@
Influence of CDIS on Convergence, Performance, and Training Efficiency. To further evaluate the influence of CDIS on convergence, the training behaviors of the baseline and baseline with CDIS are compared on the University-1652 dataset in setting i. As shown in Fig. 6(a), it presents the loss curves and their statistical values. The baseline and baseline with CDIS exhibit similar descending trends and stable convergence regions. The baseline with CDIS begins with a slightly higher loss but reaches a lower final value, which demonstrates that the optimization process remains stable and that CDIS guides the model toward a better solution. Fig. 6(b) reports the performance and the training time. The CDIS improves the performance in both the drone $\rightarrow$ satellite and satellite $\rightarrow$ drone scenarios compared to the baseline, and the additional training cost is small at about 35 seconds. Empirical evidence confirms that CDIS simultaneously enhances structural and spatial invariance, which maintains rapid convergence rates and incurs only marginal computational overhead, ultimately achieving superior performance.
Figure 7: Similarity Distribution and Feature Space Visualization on University-1652 . (a–e) show the similarity distribution visualization. The red-shaded area indicates the overlapping region between the positive and negative distributions. (f-j) show the distribution of feature embeddings in the 2D shared feature space, where circles and triangles in different colors denote drone-view and satellite-view. A total of 20 localizations are randomly selected from the test set.
@@
i ConvNext-Tiny[7] 100 60.23 81.94 - - - - - -
Sample4Geo[9] 100 80.57 96.53 80.57 96.53 86.87 98.71 91.38 99.74
DAC[48] 100 84.47 96.53 80.92 97.04 85.32 98.71 85.46 99.10
@@ IV-F Visualization
As shown in Fig. 7, to further demonstrate the effectiveness of CDIKTNet, we visualize the similarity distribution using t-SNE [44] and feature space projection.
Similarity Distribution Visualization. As shown in Fig. 7(a–b), Sample4Geo and DAC achieve overlap areas of 0.1232 and 0.1124 under limited supervision, due to their inability to exploit unpaired data. Fig. 7(c) shows that EM-CVGL [24] achieves an overlap of 0.1158. Although it performs slightly better, feature confusion remains an issue. In contrast, CDIKTNet in Fig. 7(d) achieves a lower overlap of 0.0661 with only 2% of paired data. This confirms its strong ability to align and separate cross-view features under limited supervision. Notably, in Fig. 7(e), CDIKTNet is trained in the cross-domain initialization setting described in Subsection IV-C. It still achieves an overlap of 0.0808 to outperform all methods. This shows its effectiveness in exploiting unpaired data and transferring knowledge from an external source domain.
Feature Space Visualization. In the feature space, features from the same location should cluster, while different locations should be well separated. As shown in Fig. 7(f–h), the compared methods show noticeable cross-view confusion (red circles) and dispersed feature distributions. In contrast, CDIKTNet (Fig. 7(i–j)) produces more compact and structured feature distributions in both scenarios. Although the 0% of paired data in cross-domain initialization setting shows minor confusion, the features remain more compact and more evenly aligned. This presents the ability of CDIKTNet to learn cross-view information.
Attention Heatmap Visualization. As a further demonstration of the effectiveness of SIEL, we visualize its attention heat maps in Fig. 8. Despite the large spatial differences between satellite-view and drone-view images and the feature distribution shifts caused by varying drone viewpoints, SIEL consistently highlights semantically meaningful regions in both views.